NOAA Office of Satellite and Product Operations

Please Note:  

To view imagery from the operational GOES East (GOES-16) and GOES West (GOES-17) satellites, users may visit https://www.star.nesdis.noaa.gov/goes/.

NOAA 18 HIRS Subsystem Summary

Component Description Status
CHANNEL 01Spectral Channel #01, 14.95 micro metersOrange
CHANNEL 02Spectral Channel #02, 14.71 micro metersOrange
CHANNEL 03Spectral Channel #03, 14.49 micro metersOrange
CHANNEL 04Spectral Channel #04, 14.22 micro metersOrange
CHANNEL 05Spectral Channel #05, 13.97 micro metersOrange
CHANNEL 06Spectral Channel #06, 13.64 micro metersOrange
CHANNEL 07Spectral Channel #07, 13.35 micro metersOrange
CHANNEL 08Spectral Channel #08, 11.11 micro metersOrange
CHANNEL 09Spectral Channel #09, 9.71 micro metersOrange
CHANNEL 10Spectral Channel #10, 12.47 micro metersOrange
CHANNEL 11Spectral Channel #11, 7.33 micro metersGreen
CHANNEL 12Spectral Channel #12, 6.52 micro metersGreen
CHANNEL 13Spectral Channel #13, 4.57 micro metersGreen
CHANNEL 14Spectral Channel #14, 4.52 micro metersGreen
CHANNEL 15Spectral Channel #15, 4.47 micro metersGreen
CHANNEL 16Spectral Channel #16, 4.45 micro metersGreen
CHANNEL 17Spectral Channel #17, 4.13 micro metersGreen
CHANNEL 18Spectral Channel #18, 4.00 micro metersGreen
CHANNEL 19Spectral Channel #19, 3.76 micro metersGreen
CHANNEL 20Spectral Channel #20, 0.69 micro metersGreen
ELECTRONICSInstrument ElectronicsGreen
FILTER WHEELInstrument Filter Wheel MotorsGreen
SCAN MOTORInstrument Scan MotorGreen


Subsystem History:

Date Time Component Description
03/24/201512:00:00FILTER WHEELFilter Wheel Motor had a spike in current about 257.2mA to which shortly returned back to its nominal operating value. Radiometrics saw slight change in pattern with the current spike. There were a few spikes JDAY 80 and JDAY 79 but the current value reached was less.
01/08/201423:15:00FILTER WHEELFilter Wheel Motor had a spike in current which shortly returned back to its nominal operating value. Radiometrics saw slight change in pattern with the current spike.
02/05/200912:00:00 HIRS data degraded & unsuable for the user community. Engineering health & safety is nominal
02/05/2009CHANNEL 01NOAA-18 HIRS increased IR Channel Noise
02/05/2009CHANNEL 02NOAA-18 HIRS increased IR Channel Noise
02/05/2009CHANNEL 03NOAA-18 HIRS increased IR Channel Noise
02/05/2009CHANNEL 04NOAA-18 HIRS increased IR Channel Noise
02/05/2009CHANNEL 05NOAA-18 HIRS increased IR Channel Noise
02/05/2009CHANNEL 06NOAA-18 HIRS increased IR Channel Noise
02/05/2009CHANNEL 07NOAA-18 HIRS increased IR Channel Noise
02/05/2009CHANNEL 08NOAA-18 HIRS increased IR Channel Noise
02/05/2009CHANNEL 09NOAA-18 HIRS increased IR Channel Noise
02/05/2009CHANNEL 10NOAA-18 HIRS increased IR Channel Noise
01/13/200912:00:00 NOAA-18 HIRS increased IR Channel Noise
05/10/200714:58:00SCAN MOTORHIRS to full scan mode after test completion
05/10/200713:10:00SCAN MOTORPark HIRS at space view for test
09/28/200613:05:00 Command Position Disable" return to nomral scanning. End lens test #2"
09/27/200613:15:00 Lens stabikization test #2: Commanded Position Disable then Nadir Position Disable then Nadir Position then Position Disable then IWT Position and then FW housing heater off.
09/25/200613:35:00 Commanded Scan mirror to IWT position and turn ON Filter wheel housing heater.
09/14/200618:35:00 Command Position Disable, return HIRS mirror to normal scan mode. End HIRS lens stabilization test.
09/13/200620:35:00 Command Filter Wheel Housing Heater off
09/11/200617:25:00 Command HIRS Filter Housing Heater on (for lens stabilization test #1).
09/11/200616:00:00 Command the HIRS mirror to IWT position (for lens stabilization test #1).
04/11/200606:10:00CHANNEL 01HIRS long wave channels again degraded with negative impacts on imagery.
04/07/200622:30:00CHANNEL 01HIRS long wave channels improved back to nominal values, imagery improved.
03/15/200623:11:00 Return to normal Dwell Channel
03/15/200623:10:00 Return HIRS to Full Scan.
03/15/200617:56:00 Put HIRS in Space View for 3 orbits.
03/15/200617:55:00 TIP Dwell on NHIRSMTI (HIRS Scan Mtr Current), CH245
03/14/200612:00:00CHANNEL 01HIRS payload again downgraded to YELLOW status because of long wave channel noise which has resulted in Products being unable to utilize this payload.
02/05/200604:48:00CHANNEL 01HIRS long wave channel noise continues to occur intermittently sometimes lasting for several days even in the absence of solar array moves and other potential perturbations to the loose LWIR lens cover. Channel-1 most affected.
01/04/200612:00:00CHANNEL 01Because of improvements in the long wave channel noise situation and an understanding of its causes (loose lens cover), N18 HIRS has been upgraded to Green status.
12/15/2005CHANNEL 01TOAR board investigation concluded on HIRS long wave channel noise. Problem attributed to loose LWIR lens cover that is highly sensitive to minor spacecraft attitude purturbations such as solar array offset changes. Channel-1 reported as nominal although channels 2, 4, 5, 6, 7, 9 and 10 are reported as somewhat out of spec.
10/13/200515:41:00 Return to normal Dwell Channel (bus voltage)
10/13/200515:40:00 Return HIRS to Full Scan
10/13/200514:01:00 Put HIRS in Space View
10/13/200514:00:00 TIP Dwell on NHIRFMTI (HIRS FW Mtr Current), CH261
08/15/200512:00:00CHANNEL 01HIRS payload and channel-1 component downgraded to YELLOW status because of degraded imagery (ref: TOAR 448) first detected on 6 June 05.
07/12/200517:44:00FILTER WHEELHIRS FW to low power mode
07/12/200516:21:00SCAN MOTORHIRS to full scan mode
07/12/200514:40:00FILTER WHEELHIRS FW motor to High Power
07/12/200512:57:00 Park HIRS in Space View
07/05/200518:55:00SCAN MOTORHIRS to Full Scan Mode
07/05/200517:22:00SCAN MOTORPark HIRS in Space View
06/27/200518:40:00SCAN MOTORHIRS to full scan mode
06/27/200517:00:00SCAN MOTORPark HIRS at space view
06/21/200515:00:00SCAN MOTORPlace HIRS back in full scan mode
06/21/200513:10:00SCAN MOTORPark HIRS at cold (space) view target
06/20/200523:25:00SCAN MOTORPlace HIRS in full scan mode
06/20/200521:40:00SCAN MOTORPark HIRS at warm target
06/10/200512:00:00 HIRS imagery less than nominal on several channels, under investigation.
06/09/200515:15:00SCAN MOTORHIRS to full scan mode
06/09/200513:35:00SCAN MOTORHIRS scanner park at Space View
06/08/200522:01:00SCAN MOTORHIRS to full scan mode.
06/08/200520:21:00SCAN MOTORHIRS scanner park at warm target
06/05/200515:55:00CHANNEL 20On orbit, operational. End HIRS OV period. Subsytem configured for nominal operations.
06/05/200515:54:00CHANNEL 19On orbit, operational
06/05/200515:53:55CHANNEL 18On orbit, operational
06/05/200515:53:45CHANNEL 17On orbit, operational
06/05/200515:53:35CHANNEL 16On orbit, operational
06/05/200515:53:25CHANNEL 15On orbit, operational
06/05/200515:53:15CHANNEL 13On orbit, operational
06/05/200515:53:15CHANNEL 14On orbit, operational
06/05/200515:53:10CHANNEL 12On orbit, operational
06/05/200515:53:05CHANNEL 11On orbit, operational
06/05/200515:53:00CHANNEL 10On orbit, operational
06/05/200515:52:55CHANNEL 09On orbit, operational
06/05/200515:52:50CHANNEL 07On orbit, operational
06/05/200515:52:50CHANNEL 08On orbit, operational
06/05/200515:52:45CHANNEL 06On orbit, operational
06/05/200515:52:35CHANNEL 05On orbit, operational
06/05/200515:52:30CHANNEL 04On orbit, operational
06/05/200515:52:25CHANNEL 03On orbit, operational
06/05/200515:52:20CHANNEL 02On orbit, operational
06/05/200515:52:10CHANNEL 01On orbit, operational
06/05/200515:52:00ELECTRONICSHIRS electronics turnon, filter motor to normal, calibration on.
06/04/200519:12:00 HIRS outgasssing complete. Enabled patch control heater and opened cooler cover door.
05/20/200510:22:00FILTER WHEELSpacecraft launched with HIRS filter motor powered up. HIRS detectors not yet activated, no imagery available until outgassing completes (NET 5 June)
05/20/200510:22:00SCAN MOTORSpacecraft launched with HIRS scan motor powered up. HIRS detectors not yet activated, no imagery available until outgassing completes (NET 5 June)